# Quality-Driven Curation of Remote Sensing Vision-Language Data via Learned Scoring Models

### v1

@@ 2.1.2 Vision Instruction Preference Dataset
Table 1: Comparison of different finetuned CLIP models on classification tasks. Top-1 (@1) and top-5 (@5) classification accuracies are reported. Results are evaluated on official test sets when available; otherwise, the entire dataset is used for evaluation
@@
Table 2: Comparison of different finetuned CLIP models on cross-modal retrieval tasks. We report text-to-image (T2I) and image-to-text (I2T) performance using top-1 recall (R@1) and top-5 recall (R@5)
2.2 Training and Application
Training
Our ScoreRS model is initialized with Qwen2VL-7B, with the language head replaced by a linear layer to output a scalar score, following the standard value-head-based reward model [37, 7]. We do not consider generative score modeling due to efficiency concerns. Similar to the standard recipe for training large VLMs [38, 20], we implement a multi-stage training procedure to gradually train ScoreRS to distinguish better vision-language pair. In the first stage, we train the newly introduced value head using a pure text preference dataset, UltraFeedback [6], to provide a good initialization for the value head. Then, we unfreeze the ViT and value head and train ScoreRS on our image-caption preference dataset to enable ScoreRS to better understand RS images. Finally, we unfreeze the LLM and conduct full-parameter training with our vision instruction preference dataset and the additional RLHF-V [46] dataset. This training approach enables ScoreRS to effectively identify high-quality outputs and assign higher scores to better responses across diverse RS scenarios.
Application
Beyond basic data selection using ScoreRS, we also explore its usage for RL training and BoN selection. We primarily discuss implementing ScoreRS for RL, as its applications to data selection and BoN selection are straightforward. Our RL framework is based on group relative policy optimization (GRPO) [33], chosen for its computational efficiency and ease of hyperparameter tuning. RL training methods typically require a reward model to evaluate each action trajectory and update the model parameters to favor outputs with higher scores (i.e., responses more aligned with human preferences). While DeepSeek-R1 [11] demonstrated that rule-based rewards for close-ended questions with verifiable answers are effective for RL training, this approach is insufficient for RS applications. In the RS image understanding domain, most questions are open-ended, such as ’Describe the urban development patterns visible in this satellite imagery,’ requiring nuanced interpretations rather than definitive answers. These questions also usually do not have verifiable answers, as different interpretations can lead to multiple acceptable responses. To address this challenge, we introduce a novel reward computation method that leverages ScoreRS for evaluating open-ended responses while incorporating rule-based rewards for close-ended questions, creating an approach that handles both question types. Specially, for close-ended questions, we employ a binary reward system (0 or 1) based on exact match or intersection over union (IoU) with the ground-truth answer. For open-ended questions, where we have a reference answer $y$ (typically sampled from standard vision instruction datasets), we compute the reward $r$ for a predicted answer $\hat{y}$ using the following formulation:
where $\beta>0$ is a hyperparameter controlling the reward sharpness. We use this reference-based approach because it accelerates learning. The reason behind this approach is that the reference answer serves as a baseline, allowing the policy model to improve upon it by generating responses that receive higher scores. This method focuses on continuous improvement rather than selecting the least problematic option from multiple suboptimal alternatives that may be worse than the reference answer itself. We given our detailed unsuccessful attempts and our reasoning in Sec. 10.
3 Experiment
We conduct experiments to evaluate the effectiveness of ScoreRS across three key applications: (1) vision-language data selection for training VLMs, (2) RL training, and (3) BoN selector. Additionally, we analyze the impact of our proposed multi-stage training strategy and examine the quality of our curated preference dataset.
3.1 Vision-Language Data Selection
3.1.1 CLIP Finetuning
@@
To validate the effectiveness of ScoreRS for data selection, we apply our model to score and rank the training samples from RemoteCLIP [19], which are then used to finetune CLIP-ViT-L/14 [29]. We compare our approach against three baselines: (1) the CLIP model without finetuning, (2) CLIP finetuned on the complete, unfiltered dataset, and (3) CLIP finetuned on data filtered using CLIP-score. Detailed training hyperparameters are provided in Sec. 8.2.
Figure 3: Classification and retrieval result using different percentages of data selected by CLIP-Score and ScoreRS. Results are reported as average scores across all evaluation datasets
@@
We evaluate the finetuned models on both classification and retrieval tasks. For classification, we employ a fully zero-shot setting, ensuring that none of the test set categories appear in the finetuning datasets. The results, presented in Tab. 1 and Tab. 2, challenge the conventional wisdom of ”more data yields better performance” in the context of RS-specific VLMs. Simple filtering using CLIP-score already yields improvements of 5% and 1% on classification and retrieval tasks, respectively. Our proposed ScoreRS filtering method further advances these improvements, achieving gains of 7% in classification and 2% in retrieval compared to training with the complete dataset, outperforming the CLIP-score-based filtering approach. These results support our initial hypothesis that current RS vision-language data are suboptimal, highlighting the necessity for automated quality control strategies.
We further evaluate different parameter-wise data selection methods by varying the selection thresholds. The results are presented in Fig. 3. Our analysis reveals distinct patterns for classification and retrieval tasks: retrieval tasks show higher sensitivity to noisy data, demonstrating more significant improvements with stricter filtering criteria, while classification tasks maintain reasonable performance even with relaxed selection criteria. Notably, quality filtering consistently improves model performance, aligning with previous research that highlights the crucial role of high-quality data in CLIP-style vision-language pretraining [44].
@@
AID METERML NWPU SIRI-WHU WHU-RS19 Avg. HR-Comp. HR-Pres. LR-Comp. LR-Pres. LR-R-U Avg. VG-DOIR LHRS-Bench
@@
Table 3: Comparison of different data filtering methods and selection strategies. We report accuracy on the test sets across all datasets. CLIP and ScoreRS denote data selection using CLIP score and our ScoreRS, respectively. PX indicates the selection of top Xof pretraining data in the first stage, while SX represents the selection of top Xof vision instruction data in the second stage
@@
AID METERML NWPU SIRI-WHU WHU-RS19 Avg. HR-Comp. HR-Pres. LR-Comp. LR-Pres. LR-R-U Avg. VG-DOIR LHRS-Bench
@@
Table 4: Performance comparison between our finetuned model (Qwen2VL-7B-RS) and existing vision-language models. Our model is trained on quality-filtered data comprising the top 30% of pretraining samples and 60% of vision instruction samples from the original VHM datasets, selected using ScoreRS as the quality assessment model
3.1.2 Large VLMs Finetuning
Experimental Setting
We finetune Qwen2VL-7B [38] using RS pretrained data and vision instruction datasets from VHM [27]. Following standard practices in large VLM training, we employ a two-stage finetuning approach: first training only the vision-language bridge layer using pretrained data, then training the underlying LLM using the vision instruction dataset. We investigate the effectiveness of ScoreRS in selecting high-quality data for both stages. As baselines, we compare our approach against models trained on the complete dataset and the data filtered using LongCLIP-L [48]. We opt for LongCLIP over the original CLIP model due to its ability to handle longer captions and conversations that are common in the chosen dataset. To reduce computational costs, we implement LoRA [13] for LLM finetuning in the second stage, using a rank of 8 unless otherwise specified. Detailed experimental settings and hyperparameters are provided in Sec. 8.3.
Main Result
We evaluate the finetuned model across multiple RS tasks, including image classification, visual grounding, and visual question answering. Additionally, we assess the model’s general RS knowledge using the more challenging LHRS-Bench benchmark [24]. The results are presented in  Tab. 3. For the image-caption alignment dataset, utilizing ScoreRS to filter out low-quality data and retaining only the top 30% achieves the best performance, yielding a 1% improvement on the challenging LHRS-Bench. Given that the VHM image-caption dataset is synthetically generated by Gemini-Flash, this high ranking threshold further supports our assertion regarding the importance of data quality control for RS synthetic data. We further explore the application of ScoreRS for filtering vision instruction data. Since the VHM vision instruction dataset is primarily constructed from various standard RS benchmark datasets, we find that applying an overly aggressive filtering ratio (e.g., 30%) performs slightly inferior to using the complete vision instruction dataset. However, when adopting a more moderate filtering approach (e.g., 60%), we observe substantial improvements: 3% on classification tasks, 5% on vision question answering, and 0.36% on challenging grounding and LHRS-Bench benchmarks. Notably, our ScoreRS quality assessment model consistently demonstrates superior results compared to CLIP score-based selection methods across all evaluation scenarios.
Based on these empirical findings, we strategically select the highest-scoring 30% of pretraining data and 60% of vision instruction data as ranked by ScoreRS, while scaling the LoRA rank size to 128. We then evaluate our finetuned model against leading RS-specific VLMs and state-of-the-art general-purpose vision-language models. As shown in Tab. 4, our model, trained on high-quality data selected by ScoreRS, not only achieves comparable or superior performance on classification, visual question answering, and visual grounding tasks compared to RS-specific VLMs, but also outperforms general-purpose models on the LHRS-Bench benchmark. These results further confirm that data quality in RS vision-language pairs represents the primary bottleneck limiting the full potential of VLMs for RS image understanding.
@@ 3.2 Reinforcement Learning
Building upon our fine-tuned Qwen2VL-7B-RS model, we explore the application of ScoreRS as a reward model in RL training. We randomly select 8K samples (4K open-ended and 4K close-ended) from the filtered VHM vision instruction dataset for the RL training process. Inspired by the success of Deepseek-R1 [11], we prompt Qwen2VL-7B-RS to engage in a two-step reasoning process: first thinking about the question, then providing an answer. This process is structured using special tokens: <think>, </think>for the reasoning step, and <answer>, </answer>for the response. For close-ended questions, we implement a binary reward function that assigns 1 for correct answers and 0 otherwise. For open-ended questions, we utilize ScoreRS for reward calculation as described in Sec. 2.2. Additionally, we implement a format reward to encourage proper prediction formatting. We evaluate performance on the more challenging vision grounding and LHRS-Bench tasks, as our manual analysis indicates that simpler classification and RSVQA tasks inadequately assess VLMs’ true capabilities in RS image understanding. We establish two baselines: the non-RL-trained Qwen2VL-7B-RS and an RL-trained version using only close-ended answer rewards (without ScoreRS). Our detailed settings are provided in Sec. 8.4.
The results are presented in Tab. 5. Direct application of RL training to Qwen2VL-7B-RS (i.e., Qwen2VL-7B-RS-Zero) yields performance improvements of 0.20% on LHRS-Bench and 1% on VG-DOIR. Moreover, the comparison experiment demonstrates that using ScoreRS for reward calculation outperforms the variant without ScoreRS-based rewards, supporting the effectiveness of ScoreRS as a reward model.
VG-DOIR LHRS-Bench
@@
Table 5: Performance comparison on VG-DOIR and LHRS-Bench across different RL-trained models. Qwen2VL-7B-RS-Zero denotes the model obtained by directly applying RL training to Qwen2VL-7B-RS. Qwen2VL-7B-RS-SFT refers to Qwen2VL-7B-RS fine-tuned with our manually generated reasoning data. Qwen2VL-7B-RS-R1 represents the model obtained by applying RL training to Qwen2VL-7B-RS-SFT. ”W/O ScoreRS” indicates the variant trained using RL without ScoreRS-based reward calculation
During our experiments, we observed that the reasoning patterns of Qwen2VL-7B-RS-Zero were overly simplistic. To further maximize the benefits of RL training, we implemented a multi-stage approach. First, we used Qwen2VL-7B-RS-Zero to answer our collected RS-specific questions introduced in Sec. 2.1.2. We then manually reviewed and refined the reasoning responses from the generated results, creating a curated dataset of 2K manually annotated reasoning-answer pairs. Finally, we fine-tuned Qwen2VL-7B-RS using this curated dataset before applying RL training with the same training data used for Qwen2VL-7B-Zero. The resulting model, Qwen2VL-7B-RS-R1, demonstrates significant improvements with this strategy, achieving gains of over 5% on VG-DOIR and 2% on LHRS-Bench, while outperforming the variant trained without ScoreRS-based reward calculation. Our analysis indicates that the reasoning patterns have become more reasonable. Representative conversation examples are provided in Fig. 7.
@@ 3.3 Best-of-N Selection
We employ our ScoreRS model as a BoN selector. Specifically, for each question, we generate multiple candidate answers and used ScoreRS to score these question-answer pairs, ultimately selecting the highest-scoring answer for evaluation. To thoroughly evaluate the effectiveness of ScoreRS as a BoN selector, we utiliz three models for answer generation: the RS-specific LHRS-Bot-Nova, the general-purpose Qwen2VL-7B, and our fine-tuned Qwen2VL-7B-RS-R1. We conduct evaluations on both VG-DOIR and LHRS-Bench datasets. Due to computational constraints and the large volume of the VG-DOIR dataset, we sample the first 1,000 instances for evaluation to reduce experimental costs. For all the answer generation, we set the sampling temperature and top-p parameters to 0.95 and 1, respectively.
Figure 4: Performance improvements when using ScoreRS as a BoN selector. We employ ScoreRS to identify the optimal response from multiple generated candidates for each query.
The results are presented in Fig. 4, which clearly demonstrates that scaling computation at test time and employing our ScoreRS as a selector consistently improves evaluation performance across different models on both complex perception tasks (VG-DOIR) and holistic vision understanding tasks (LHRS-Bench). Notably, for the challenging LHRS-Bench dataset, using ScoreRS as the BoN selector enables accuracy to exceed 70% for the first time. This promising result further indicates the potential of ScoreRS not only for data selection but also for test-time scaling of base models in RS image understanding.
@@ 3.4 Ablation Analysis
To validate the effectiveness of our multi-stage training strategy and the quality of our preference dataset, we conduct ablation studies on a held-out evaluation set. The evaluation dataset comprises 6K samples, consisting of 2K randomly selected pairs from our image-caption preference dataset and 4K samples from the vision instruction preference dataset. We evaluate the performance using accuracy metrics, where success is defined as the scoring model assigning a higher score to the preferred text ($T^{+}$) compared to its counterpart ($T^{-}$).
We begin by validating the effectiveness of our preference dataset through joint training experiments. The joint training process consists of two stages: (1) an initial stage using pure text preferences to initialize the value head, which we find crucial for training stability, and (2) a subsequent stage where both the ViT and LLM are unfrozen. Our experiments demonstrate that training with our complete preference dataset yields the highest reward accuracy (Tab. 6). Notably, excluding any subset of our constructed dataset leads to decreased performance even when evaluating on the same domain (e.g., excluding the image-caption preference dataset results in inferior performance on the image-caption evaluation subset compared to joint training with all data). Furthermore, implementing our proposed multi-stage training strategy achieves best results, improving the reward accuracy by more than 10% compared to joint training. These results demonstrate both the effectiveness of our curated preference dataset and the clear benefits of our multi-stage training strategy.
@@
Table 6: Ablation study examining the impact of multi-stage training strategy and preference dataset composition. We first evaluate the contribution of each preference dataset component through joint training experiments, then demonstrate the benefits of our proposed multi-stage training approach compared to joint training alternatives. (P.D. = Preference Dataset, ICPD = Image-Caption Preference Dataset, VIPD = Vision Instruction Preference Dataset)
4 Related Work
4.1 RS Vision-Language Data Curation
Unlike the general vision domain, where vision-language data can be readily crawled from abundant image-text interleaved webpages [44, 9, 36, 30, 4], the RS domain presents a unique challenge. While rich in open-source image data [24, 23, 35], RS lacks corresponding text descriptions and analytical conversations about these images. To harness the potential of VLMs in the RS domain, existing studies have primarily focused on two approaches for data curation: rule-based methods and synthetic data generation using established VLMs. For rule-based approaches, RS5M [51] explores the use of binary RS image classifiers and filtering methods to select RS image-text pairs from open-source general vision-language datasets. Similarly, RemoteCLIP [19], Skyscript [39], and SkysenseGPT [22] implement manually designed text templates and relation graphs for data construction. In parallel, studies such as VHM [27], LHRS-Bot [24], LHRS-Bot-Nova [18], and GeoChat [17] leverage flagship VLMs like GPT4 or Gemini for synthetic vision-language data curation. While these works have advanced the development of RS-specific VLMs, our practical applications reveal that data quality remains suboptimal, motivating the development of automated data selection methods.
4.2 RS VLMs
The remarkable success of VLMs in understanding images and engaging in complex tasks [21, 26] has inspired the development of specialized VLMs for RS image interpretation. Several approaches have focused on adapting CLIP [29] for the RS domain. RemoteCLIP [19], Skyscript [39], and RS5M [51] fine-tuned CLIP using carefully curated RS image-caption pairs, enabling robust RS zero-shot classification and image-text retrieval capabilities. For larger VLMs, GeoChat [17] pioneered the finetuning of LLaVA [20] with RS-specific vision instruction datasets. Subsequent works have expanded this paradigm through various strategies: incorporating volunteer geographic information (VGI) [24], integrating high-quality vision-language data [18, 27], modeling relational graphs [22], implementing ensemble vision encoders for vision-centric designs [50], and improving temporal understanding of RS images [14]. The versatile and successful applications of RS VLMs have directed our attention to the fundamental building blocks underlying these models: data quality and utilization. Rather than proposing new architectural designs, our work focuses on addressing how to improve data quality and enhance VLM performance within existing frameworks.
4.3 Parameter-Wise Data Selection
As data volume continues to expand and data complexity evolves, quality control has become increasingly challenging. Beyond traditional rule-based data selection methods [34, 40], researchers have shifted focus toward parameter-wise selection approaches for more sophisticated data curation. In the language domain, methods such as QaRater [41] and LESS [43] explore training dedicated scoring models or leveraging model gradients to select high-quality data for LLM pretraining and instruction tuning. Similarly, for vision-language data, Zhang et al. [49] investigated scoring models for curating image-text pairs. The model selection approach has been widely adopted in recent state-of-the-art VLMs, including Llama-3.2-Vision [10], DeepseekVL2 [42], and Qwen2VL [38], all of which implement scoring models for data filtering. Despite these advances, there is a notable absence of open-source scoring models specifically designed to evaluate the quality of RS vision-language data. In this work, we develop and release the first open-source scoring model tailored for RS vision-language data quality assessment.

### v2

@@ 4.1.1 CLIP Finetuning
Table 2: Comparison of finetuned CLIP models on cross-modal retrieval tasks. Text-to-image (T2I) and image-to-text (I2T) performance shown using top-1 recall (R@1) and top-5 recall (R@5)
@@
Figure 3: Classification and retrieval results using different percentages of data selected by CLIP-Score and ScoreRS. Top-1 (@1) results shown as average scores across all datasets
We further evaluate different parameter-wise data selection methods across various thresholds (Figure 3). Our analysis reveals task-specific patterns: retrieval tasks show higher sensitivity to noise, with greater improvements under stricter filtering, while classification tasks maintain reasonable performance with more relaxed criteria. Notably, quality filtering consistently improves model performance, aligning with previous research that highlights the crucial role of high-quality data in CLIP-style pretraining (xu2023demystifying, ). More discussion of these results and further comparison in even more extreme data filtering scenarios can be found in Appendix F.1 and C.3.2.
4.1.2 Large VLMs Finetuning
@@
We fine-tune Qwen2VL-7B (wang2024qwen2, ) using RS image-caption data and vision instruction datasets from VHM (pang2024vhmversatilehonestvision, ). Following standard VLM training practices, we use a two-stage approach: first training only the vision-language bridge layer with pretrained data, then training the LLM with vision instructions. We evaluate ScoreRS’s effectiveness in selecting high-quality data for both stages, comparing against models trained on the complete dataset and data filtered by LongCLIP-L (zhang2024long, ), which we chose over original CLIP for its superior handling of longer captions and conversations. To reduce computational costs, we implement LoRA (hu2021lora, ) with rank 8 for LLM finetuning. Detailed configurations are provided in Appendix D.4.
@@
We evaluate our finetuned model on multiple RS tasks (image classification, visual grounding, question answering) and assess general RS knowledge using the challenging LHRS-Bench (muhtar2024lhrs, ). As shown in Table 3, filtering image-caption data with ScoreRS and retaining the top 30% achieves best performance, yielding a 1% improvement on LHRS-Bench. Given that the VHM image-caption dataset is synthetically generated by Gemini-Flash, this high ranking threshold further supports our assertion regarding the importance of quality control for RS synthetic data. We further explore the application of ScoreRS for filtering vision instruction data. Since VHM vision instruction dataset derives from standard RS benchmarks, applying aggressive filtering (e.g., 30%) performs slightly worse on classification than using the complete dataset. However, when adopting a more moderate filtering approach (e.g., 60%), we observe substantial improvements: 3% on classification tasks, 5% on vision question answering, and 0.36% on challenging grounding and LHRS-Bench benchmarks. Notably, ScoreRS consistently outperforms CLIP score-based selection across all evaluation scenarios.
Based on these empirical findings, we strategically select the highest-scoring 30% of pretraining data and 60% of vision instruction data as ranked by ScoreRS, while scaling the LoRA rank size to 128. We then evaluate our finetuned model against leading RS-specific VLMs and state-of-the-art general-purpose VLMs. As shown in Table 4, our model, trained on high-quality data selected by ScoreRS, not only achieves comparable or superior performance on classification, visual question answering, and visual grounding tasks compared to RS-specific VLMs, but also outperforms general-purpose models on the LHRS-Bench benchmark. These results further confirm that data quality in RS vision-language pairs represents the primary bottleneck limiting the full potential of VLMs for RS image understanding. More discussion of these results can be found in the Appendix F.4.
Table 3: Comparison of different data filtering methods and selection strategies. PX indicates the selection of top X% of image-caption data in the first stage, while SX represents the selection of top X% of vision instruction data in the second stage
@@
AID METERML NWPU SIRI-WHU WHU-RS19 Avg. HR-Comp. HR-Pres. LR-Comp. LR-Pres. LR-R-U Avg. VG-DIOR LHRS-Bench
@@
Table 4: Comparison between our finetuned model (Qwen2VL-7B-RS) and existing VLMs. Our model is trained on quality-filtered data comprising the top 30% of pretraining samples and 60% of instruction samples from the VHM datasets, selected using ScoreRS as the quality assessment model
@@
AID METERML NWPU SIRI-WHU WHU-RS19 Avg. HR-Comp. HR-Pres. LR-Comp. LR-Pres. LR-R-U Avg. VG-DIOR LHRS-Bench
@@ 4.2 Reinforcement Learning
We explore using ScoreRS as a reward model for RL training of our Qwen2VL-7B-RS model, using 8K samples (4K open-ended, 4K close-ended) from the filtered VHM vision instruction dataset. Inspired by Deepseek-R1 (guo2025deepseek, ), we implement a two-step reasoning process with special tokens (<think>, </think> for reasoning; <answer>, </answer> for responses). Our reward functions include binary rewards for close-ended questions, ScoreRS-based rewards for open-ended questions (Section 3.3), and format rewards for proper structuring. We evaluate on vision grounding and LHRS-Bench tasks, which better assess RS image understanding capabilities than simpler tasks, comparing against non-RL-trained Qwen2VL-7B-RS and an RL version without ScoreRS rewards (where we substitute 4K additional close-ended questions to maintain equal training steps). Detailed settings are presented in Appendix D.5.
Table 5: Comparison of RL-trained models. Qwen2VL-7B-RS-Zero: Directly apply RL to Qwen2VL-7B-RS. Qwen2VL-7B-RS-SFT: Qwen2VL-7B-RS fine-tuned with our manually generated reasoning data. Qwen2VL-7B-RS-R1: RL applied to Qwen2VL-7B-RS-SFT. "w/o ScoreRS": variant trained without ScoreRS-based rewards
VG-DIOR LHRS-Bench
@@
Table 5 shows that direct RL training on Qwen2VL-7B-RS (Qwen2VL-7B-RS-Zero) improves performance by 0.2% on LHRS-Bench and 1% on VG-DIOR. Notably, using ScoreRS for reward calculation outperforms the variant without ScoreRS-based rewards, confirming ScoreRS’s effectiveness as a reward model. We also provide a comparison of this training paradigm with DPO and compare it with using GPT-4o as a reward model for open-ended questions in Appendix C.4.
During experiments, we observe that Qwen2VL-7B-RS-Zero exhibits overly simplistic reasoning patterns. To maximize RL training benefits, we implement a multi-stage approach: first, we used Qwen2VL-7B-RS-Zero to answer our RS-specific questions from Section 3.2.2, then manually refined these responses to create 2K curated reasoning-answer pairs. We fine-tune Qwen2VL-7B-RS on this dataset before applying RL training with the same data used for Qwen2VL-7B-Zero. The resulting model, Qwen2VL-7B-RS-R1, shows significant improvements—gaining over 5% on VG-DIOR and 2% on LHRS-Bench—while also outperforming variants trained without ScoreRS-based rewards. Analysis confirms more reasonable reasoning patterns, with conversation examples in Appendix 13.
@@ 4.3 Best-of-N Selection
We compare our ScoreRS and CLIP-score with LongCLIP zhang2024long as a BoN selector Specifically, for each question, we generate multiple candidate answers and select the highest-scoring answer with given selector. We utilize three models for answer generation: the RS-specific LHRS-Bot-Nova, the general-purpose Qwen2VL-7B, and our fine-tuned Qwen2VL-7B-RS-R1. We evaluate on both VG-DIOR and LHRS-Bench datasets. Due to computational constraints and the large volume of the VG-DIOR dataset, we sample the first 1,000 instances for evaluation to reduce experimental costs. For all the answer generation, we set the sampling temperature and top-p parameters to 0.95 and 1, respectively.
Figure 4: Comparison with different BoN selectors.
Figure 4 demonstrates that using ScoreRS as a selector consistently enhances evaluation performance across models for both complex perception (VG-DIOR) and holistic vision understanding (LHRS-Bench) tasks, while CLIP-score underperforms in these scenarios. Notably, with ScoreRS as the BoN selector, accuracy on the challenging LHRS-Bench dataset exceeds 70% for the first time, highlighting ScoreRS’s potential not only for data selection but also for test-time scaling of base models in RS image understanding. We provide additional model validations and comparisons with majority voting in Appendix C.5.
@@ 4.4 Ablation Analysis
We validate our multi-stage training strategy and preference dataset quality through ablation studies on a held-out set of 6K samples (2K from image-caption preferences, 4K from vision instruction preferences). Performance is measured by accuracy, defined as the scoring model correctly assigning higher scores to preferred texts ($T^{+}$) over their counterparts ($T^{-}$). Additional ablation analyses examining score model size and the benefits of RS-specific VLM initialization are provided in the Appendix C.2.
Table 6: Ablation study on multi-stage training strategy and preference dataset composition. We evaluate each dataset component’s contribution (ICPD = Image-Caption Preference Dataset, VIPD = Vision Instruction Preference Dataset) and demonstrate our multi-stage approach’s advantages over joint training alternatives.
@@
We validate our preference dataset’s effectiveness through joint training experiments comprising two stages: (1) initializing the value head with pure text preferences for stability, and (2) unfreezing both ViT and LLM components. As shown in Table 6, training with our complete preference dataset achieves the highest reward accuracy. Notably, omitting any subset degrades performance even when evaluating on different domain (e.g., excluding image-caption preferences reduces accuracy on the vision instruction evaluation subset). Our multi-stage training strategy further improves reward accuracy by over 10% compared to joint training. These results confirm both the quality of our curated preference dataset and the benefits of our training approach.